The formula used to compute a large-sample confidence interval for p is

p̂ ± (z critical value)

p̂(1 − p̂)
n

What is the appropriate z critical value for each of the following confidence levels? (Round your answers to two decimal places.)

a) 85%
Answer #1

solution:

At 85% confidence level the z is ,

= 1 - 85% = 1 - 0.85 = 0.15

/ 2 = 0.15/ 2 = 0.075

Z/2 = Z0.075 =1.44 ( Using z table    )